# Anytime Replanning of Robot Coverage Paths for Partially Unknown Environments

### v1

@@ VII-B Coverage cost comparisons
Figure 12: Comparison of the replanning approaches using the varying clutter dataset: (a) Difference in coverage cost from an offline optimal planner (all obstacles known). (b) Minimum stop time of the robot during coverage (GD omitted as replanning time is negligible). The line shows the average and the lighter region shows the data interval.
We compare the total coverage cost (drive time $+$ stop time) of our two proposed OARP-Replan approaches against two other replanning approaches: (i) GD replan, which resembles the practical case of following a turn-minimizing coverage path with a local planner; and (ii) the OARP approach as proposed in [6]. OARP is an offline planner, so we modify it to be usable online for replanning. We refer to this online approach as OARP-Original which is implemented as follows: (i) ranks are replanned by solving the linear relaxation of MILP-0 from Section V (identical to the MILP in [6]), and (ii) the tour of the resulting ranks is computed by solving the GTSP as formulated in Section VI.
@@
We first review the results from the fixed clutter dataset. For each base environment, we ran 5 trials, where each trial was generated with a new random set of obstacles occupying 10% of the free space in the environment. Fig. 10(a) and (b) show the results of the comparison where we plot the coverage cost and the minimum stop time for each method respectively. To compare coverage costs, we plotted the cost difference from that of an offline planner (OARP) provided with all obstacle information a-priori. We observe that both OARP-Replan approaches result in more optimal coverage paths relative to GD replan, with an average improvement of about 13% across all maps. To visualize this improvement, we show an example map with three unknown obstacles (red boxes) in Figs. 11(a) and (b) where the initial path was replanned using GD replan and OARP-Replan-1 respectively. Fig. 11(c) shows two examples where we observe qualitative differences between the approaches. In example 1, paths replanned using GD replan (red paths) have multiple long detours around the obstacle, whereas OARP-Replan-1 has no such detours. In example 2, GD replan generates messy overlapping paths when avoiding the large obstacle, while OARP-Replan-1 reorients the ranks in this region to obtain a path with fewer overlaps. Using OARP-Replan-2, we observe similar improvements in the replanned path for this example. Since the tour of the replanned ranks is computed with little to no stoppage, the replanning time does not affect the total coverage cost.
@@
When comparing with OARP-Original, both OARP-Replan approaches achieve more optimal coverage, with the gap increasing as initial paths get longer (larger maps). This is because MILP-0 does not limit the change to the coverage path, leading to longer replanning times and resulting robot stoppages (Fig. 10 (b)). In contrast, OARP-Replan approaches cause little to no stoppage. We expect this difference in performance to be exaggerated in weaker computation setups, such as onboard computers on robot platforms. Between the OARP-Replan approaches, OARP-Replan-1 consistently produces slightly better path costs than OARP-Replan-2.
@@
We conducted similar experiments using the variable clutter dataset, where we ran 5 trials for each clutter level. Fig. 12 shows the results of the coverage cost comparisons. In Fig. 12(a), we observe that the cost difference between the OARP-Replan approaches and GD replan increases with increasing clutter, where OARP-Replan improves the cost by 18% on average. This is because, with more obstacles, GD replan causes many detours, while the OARP-Replan approaches tend to shorten such detours. In comparison with OARP-Original, the OARP-Replan approaches plan slightly more costly coverage paths. However, OARP-Original tends to experience longer stops with increasing clutter (Fig. 12(b)). We also observe some stoppage with using the OARP-Replan approaches in environments of 20% clutter. We believe this is because we do not update our runtime estimator during an experiment, making it inaccurate for highly cluttered environments (e.g. computing transitions take more time with more obstacles). A more sophisticated estimator that updates itself during coverage may eliminate this stoppage.
VII-C MILP runtime comparisons
@@
In this section, we compare the average runtimes of the rank replanning MILPs for both OARP-Replan approaches. Our approach relies on the MILP being solved efficiently so that the GTSP tour can be reliably computed within the time budget (preventing robot stoppages). The MILP from OARP-Original is excluded from this comparison as it does not constrain the GTSP input size.
@@
For the varying clutter dataset, we observe that the average runtimes of both MILPs are comparable, with MILP-1 performing better overall. Increasing clutter seems to slightly change both MILPs’ runtimes, and as such, we observed no drastic runtime difference as with the fixed clutter dataset.
VII-D Comparison against BCD
In addition to the above comparisons, we compared the proposed OARP-Replan (i.e. OARP-Replan-1) approach against the Boustrophedon Cell Decomposition (BCD) method from [7] (i.e., comparing an offline solution to an online solution). For this comparison, we use the dataset from the BCD paper [7] that consists of around 300 scans of environments with varying numbers of obstacles or holes. The outer boundary of each map is a square, which is then filled with the obstacles observed in the scans. The maps are arranged in the order of increasing complexity, which is given by the number of hole vertices (e.g., one square hole consists of 4 vertices). To perform this comparison, we use similar robot parameters as in [7]: a tool width of 3 $m$, maximum velocity of 3 $m/s$, linear acceleration of ±$\pm 1$ $m/s^{2}$, and turn-in-place dynamics with an angular velocity of $30^{\circ}/s$.
BCD is an offline planning method, which requires prior information about the obstacles in the environment. While there are online variations of BCD, such as $*{}^{*}$, the BCD approach as proposed in [7] aims to minimize turns and is therefore chosen as our baseline. For this comparison, we provide all obstacle information to the offline BCD planner, while providing no information to OARP-Replan. OARP-Replan must therefore replan an initial path covering the unobstructed base environment (a simple square) whenever the robot detects obstacles. We also add the results of using the offline OARP planner from [6], as it represents OARP-Replan when provided with all obstacle information and unlimited solver time. For clarity, we will refer to the two offline approaches as BCD-Offline and OARP-Offline for the rest of this section. Fig. 14 shows the results of comparing the total coverage cost of using all approaches. We observed that both OARP-Replan and OARP-Offline outperform BCD-Offline across all maps. We observed that OARP-Replan improves the coverage cost by about 18.6% on average from BCD-Offline, while OARP-Offline improves the cost by about 25.5%. We believe this improvement to be due to the optimal coverage lines produced by both OARP-based methods, as BCD-Offline paths tend to overlap closer to cell boundaries and obstacles. We conclude that using OARP-Replan with a simple initial path is a competitive alternative to using BCD-Offline with complete obstacle information. However, if obstacle information is available beforehand, OARP-Offline is a better alternative as it computes truly optimal paths.
@@ VII-D Comparison against BCD
Figure 13: Comparison of average MILP runtimes for the OARP-Replan approaches. (a) Fixed clutter dataset comparisons, with maps sorted by base map area. (b) Varying clutter dataset comparisons. The line shows the average and the lighter region shows the data interval.
VII-E ROS simulation case study
@@
We conclude this section by providing a case study of a Turtlebot 3 replanning coverage paths in a ROS simulation using GD replan and OARP-Replan (i.e. OARP-Replan-1). This is to demonstrate the real-world applicability of our replanning approach. We generated an anonymized environment as shown in Fig. 15, for which we generated an initial coverage path as shown in Fig. 15(a). We then added new obstacles to the environment that occupied 10% of the initial area. As the robot observes new obstacles blocking this path, the robot replans the path either using OARP-Replan or GD replan. The replanning is done while driving and the robot switches to the new path when available. The robot is also equipped with a local planner to traverse the path and make small changes if the path goes close to the walls. Fig. 15(b) shows a snapshot of the robot replanning the coverage path using OARP-Replan. We use the robot parameters from Table I for this simulation with some changes to use the default Turtlebot controller in ROS: we set maximum velocity to 0.3 $m/s$ and linear acceleration to 1 $m/s^{2}$. With these parameters, we observed that OARP-Replan covered the environment 16.5% more optimal than GD replan along a path that was 12.2% shorter. This difference is mainly due to the reduction in long detours where the robot travels around obstacles many times. We included recordings of this case study in our video attachment.
Figure 14: Coverage cost comparisons between OARP-Replan, OARP-Offline from [6], and BCD-Offline from [7]. The offline methods are given all obstacle information beforehand, while OARP-Replan has no obstacle information and replans an initial path online.
@@
VIII Conclusion
@@
We proposed an anytime coverage replanning approach called OARP-Replan to replan coverage paths for environments with unknown static obstacles. OARP-Replan solves this problem in two stages: (i) rank replanning, and (ii) touring. The first stage replans the ranks (straight-line paths) of an obstructed coverage path, while the second stage computes a tour of the replanned ranks to obtain the new path. We proposed two mixed-integer linear programs (MILPs) that each solve the rank replanning stage so that the resulting touring stage is completed within a time budget $\tau$. Simulations on maps of real-world environments showed that the OARP-Replan variants using each MILP achieve better coverage performance than both a greedy replanner and a state-of-the-art coverage planner modified for online replanning [6]. Moreover, we also compared OARP-Replan with an offline approach (treating all obstacles as known) [7] where we observed a 19% improvement in coverage cost.
@@
For future work, we look to extend this approach to address dynamic obstacles. This would involve detecting areas that could be free at a later time and coming back to cover them without long detours. Another future direction involves replanning proactively using predictions about environment occupancy to balance coverage and exploration, which can be competing objectives. Removing the axis-parallel constraint and allowing multiple directions of coverage may also improve coverage paths for non-orthogonal environments.

### v2

@@ VII-B Comparisons with online approaches
Fixed Clutter Dataset: We first review the results for the fixed clutter dataset. For each base environment, we ran 5 trials, where each trial was generated with a new random set of obstacles occupying 10% of the free space in the environment. The maps are sorted in the order of base coverage time, which is the time to cover the base environment (without the unknown obstacles) using the initial coverage path. Fig. 11(a) and (b) show the results of the comparison where we plot the following for each method: (a) the difference between the total coverage time (including stops) and the base coverage time (i.e. total -- base), and (b) the minimum stop time. To visualize the best-case coverage time for each environment (a lower bound), we also include the results for an offline planner (OARP) when provided with all obstacle information a-priori. We observe that both OARP-Replan approaches result in lower total coverage times relative to all baseline approaches, and comes closest to the performance of the offline planner. In comparison with GD replan, OARP-Replan exhibits an average improvement of about 12% across all maps. To visualize this improvement, we show an example map with three unknown obstacles (red boxes) in Figs. 12(a) and (b) where the initial path was replanned using GD replan and OARP-Replan-1 respectively. Fig. 12(c) shows two examples where we observe qualitative differences between the approaches. In example 1, paths replanned using GD replan (red paths) have multiple long detours around the obstacle, whereas OARP-Replan-1 has no such detours. In example 2, GD replan generates messy overlapping paths when avoiding the large obstacle, while OARP-Replan-1 reorients the ranks in this region to obtain a path with fewer overlaps. Using OARP-Replan-2, we observe similar improvements in the replanned path for this example. Since the tour of the replanned ranks is computed with little to no stoppage, the replanning time does not affect the total coverage time.
@@
When comparing with OARP-Original and the HR approach [18], both OARP-Replan approaches achieve faster coverage times, with the gap increasing as initial paths get longer (larger maps). This is because both ILP-0 and the HR approach do not limit the change to the coverage path, leading to longer replanning times and resulting robot stoppages (Fig. 11 (b)). In contrast, OARP-Replan approaches cause little to no stoppage. We expect this difference in performance to be exaggerated in weaker computation setups, such as onboard computers on robot platforms. The OARP-Replan approaches also outperform the SB approach as a result of minimizing the number of ranks (turns) in the coverage path. However, we observe that the SB approach performs well on the map with the longest base coverage time when compared to the other baselines. Note that this map is not the largest in terms of area, so we reason that the performance of SB is more closely related to map area than to base coverage time (which may be influenced by complex layouts/geometry). Between the OARP-Replan approaches, OARP-Replan-1 consistently plans paths with slightly shorter coverage times than OARP-Replan-2.
Figure 13: Comparison of the online replanning approaches using the varying clutter dataset: (a) Difference between total coverage time and base coverage time. (b) Minimum stop time of the robot during coverage (GD and SB omitted as replanning time is negligible). The line shows the average and the lighter region shows the data interval.
@@
Varying Clutter Dataset: We also conducted similar experiments using the varying clutter dataset, where we ran 5 trials for each clutter level. The map chosen for this dataset has a base coverage time of 944s. Fig. 13 shows the results of the coverage time comparisons. In Fig. 13(a), we plot the difference between the total coverage time and base coverage time. We observe that the OARP-Replan approaches outperform GD replan by 18% on average, with the performance gap increasing with clutter level. This is because, with more obstacles, GD replan causes many detours, while the OARP-Replan approaches tend to shorten such detours.
In comparison with OARP-Original and the HR approach [18], the OARP-Replan approaches plan coverage paths with comparable coverage times. However, both OARP-Original and HR experience longer stops with increasing clutter (Fig. 13(b)). We also observe some stoppage with using the OARP-Replan approaches in environments of 20% clutter. We believe this is because we do not update our runtime estimator during an experiment, making it inaccurate for highly cluttered environments (e.g. computing transitions take more time with more obstacles). A more sophisticated estimator that updates itself during coverage may eliminate this stoppage. We also observe that the coverage time using the SB approach plateaus with increasing clutter due to SB’s rapid online computation of coverage paths. Although OARP-Replan outperforms SB at all the clutter levels tested, this performance gap may decrease with higher clutter levels ($>20\%$).
VII-C ILP runtime comparisons
@@
In this section, we compare the average runtimes of the rank replanning ILPs for both OARP-Replan approaches. Our approach relies on the ILP being solved efficiently so that the GTSP tour can be reliably computed within the time budget (preventing robot stoppages). The ILP from OARP-Original is excluded from this comparison as it does not constrain the GTSP input size.
@@
For the varying clutter dataset, we observe that the average runtimes of both ILPs are comparable, with ILP-1 performing better overall. Increasing clutter seems to slightly change both ILPs’ runtimes, and as such, we observed no drastic runtime difference as with the fixed clutter dataset. While this comparison was performed with a fixed tool width (grid cell size), changing the tool width could either increase or decrease the number of grid cells for all maps. However, since the change in map area also has the same effect, we expect the observed trend to be maintained with a change in tool width.
@@
Figure 14: Comparison of average ILP runtimes for the OARP-Replan approaches. (a) Fixed clutter dataset comparisons, with maps sorted by base map area. (b) Varying clutter dataset comparisons. The line shows the average and the lighter region shows the data interval.
VII-D Comparisons with offline approaches
In this section, we compare the proposed OARP-Replan approach (i.e. OARP-Replan-1) against offline coverage planning approaches, which are provided with all obstacle information prior to coverage. In contrast, we provide OARP-Replan with no prior obstacle information, so it must replan the coverage path whenever the robot detects obstacles. We compare OARP-Replan against 3 offline approaches: (i) boustrophedon cell decomposition (BCD) method [7], (ii) line coverage (LC) approach [8], and (iii) the offline OARP planner [6]. For clarity, we will refer to the offline approaches as BCD-Offline, LC-Offline, and OARP-Offline for the rest of this section. The BCD-Offline algorithm also resembles the BCD approach from the ‘ipa_room_exploration’ ROS package, which is the most effective algorithm in the package according to the adjoining survey in [9].
While OARP-Replan was developed with mobile robots as a focus, it can also be used for coverage planning with aerial robots (fixed coverage tool width). As such, we apply the OARP-Replan approach to plan coverage paths for an aerial robot where the obstacles in the environment are not previously known. We use the dataset from the BCD paper [7] that consists of around 300 scans of environments with varying numbers of obstacles or holes. The outer boundary of each map is a square, which is then filled with the obstacles observed in the scans. The maps are arranged in the order of increasing complexity, which is given by the number of hole vertices (e.g., one square hole consists of 4 vertices). We also use robot parameters similar to that of the UAV from [7]: a tool width of 3 $m$, maximum velocity of 3 $m/s$, and linear acceleration of ±$\pm 1$ $m/s^{2}$ (turning time omitted to match [7]). Additionally, OARP-Replan is provided with an initial path for the unobstructed base environment (a square), and a radial range sensor similar to previous sections to identify obstacles.
Figure 15: Comparison results with offline planning approaches. (a) Coverage time comparisons between OARP-Replan and offline baselines. (b) Violin plot of percent coverage achieved by the OARP-based approaches (with data points plotted). BCD-Offline and LC-Offline are omitted as they are exact approaches that achieve 100% coverage.
Fig. 15(a) shows the results of comparing the total coverage times of using all approaches. We observe that OARP-Replan outperforms BCD-Offline and LC-Offline across all maps, with an improvement of about 25% on average from BCD-Offline. We believe this improvement to be due to the optimal coverage lines produced by both OARP-based methods, as BCD-Offline paths tend to overlap closer to cell boundaries and obstacles. We also observe that OARP-Offline exhibits the best performance with an average improvement of 31% in coverage time from BCD-Offline. However, note that OARP-Offline is equivalent to OARP-Replan when provided with all obstacle information a-priori and unlimited solver time. We conclude that using OARP-Replan with a simple initial path is a competitive alternative to using BCD-Offline with complete obstacle information. However, if obstacle information is available beforehand, OARP-Offline is a better alternative as it computes paths with the fastest coverage times.
One thing to note is that both BCD-Offline and LC-Offline are exact approaches that aim to cover the entire environment (100% coverage). In contrast, the OARP-based approaches use grid approximations that may not cover the environment fully. To address this, we plot the percent coverage achieved by both OARP-based methods (they cover the same grid) in Fig. 15(b), where we observe that the OARP-based approaches achieve about 96% coverage on average across all environments, with missed coverage usually occurring close to obstacles and boundaries. Therefore, we recommend OARP-Replan for covering environments with unknown obstacles where small amounts of missed coverage is acceptable.
VII-E Simulation case study
@@
This section provides a case study of a Turtlebot 3 replanning coverage paths in a ROS simulation using GD replan and OARP-Replan (i.e. OARP-Replan-1). This is to demonstrate the real-world applicability of our replanning approach in contrast to GD replanning. We generated an anonymized environment as shown in Fig. 16, for which an initial coverage path was planned (Fig. 16(a)). We added new obstacles to the environment that occupied 10% of the initial area. As the robot observes new obstacles blocking this path, the robot replans the path either using OARP-Replan or GD replan. The replanning is done while driving and the robot switches to the new path when available. The robot is also equipped with a local planner to traverse the path and make small changes if the path goes close to the walls. Fig. 16(b) shows a snapshot of the robot replanning the coverage path using OARP-Replan. We use the robot parameters from Table I for this simulation with some changes to use the default Turtlebot controller in ROS: we set maximum velocity to 0.3 $m/s$ and linear acceleration to 1 $m/s^{2}$. With these parameters, we observed that OARP-Replan covered the environment 16.5% faster than GD replan along a path that was 12.2% shorter. This difference is mainly due to the reduction in long detours where the robot travels around obstacles many times. We included recordings of this case study in our video attachment.
@@
VIII Robot Experiments
In this section, we present the results of testing OARP-Replan (i.e. OARP-Replan-1) on an Avidbots Neo robot (Fig. 17(a)) covering a partially unknown environment. The robot’s task is to cover a test environment for which an initial path is provided (Fig. 18(a)). The robot has tricycle dynamics and is equipped with a local planner to make small changes to difficult parts of the initial path. However, the robot may encounter large obstacles during coverage (Fig. 17(b)) which may result in narrow regions that the robot cannot reach along its current path. Therefore, a replan of the coverage path is necessary to reach such regions and maximize the coverage of the environment.
We conducted our experiments using two test scenarios. Fig. 18 shows the results of applying OARP-Replan in both scenarios. In scenario 1 (Fig. 17(a) / Fig. 18(b)), the robot observes an obstacle that interrupts a large number of ranks in the initial path. The path is replanned using OARP-Replan, which computes horizontal ranks to comfortably cover the narrow region (bottom right of the environment in Fig. 18(b)). In scenario 2 (Fig. 17(b) / Fig. 18(c)), the robot encounters a large obstacle that it must observe over time. Later in the task, the robot observes a part of the obstacle that causes a narrow coverage region (top right of the environment in Fig. 18(c)). OARP-Replan allows the robot to cover this region by re-orienting the coverage ranks. We provide recordings of these experiments in our multimedia attachment.
To implement OARP-Replan, we require an occupancy map that uses the LIDAR scan collected by the robot to identify grid cells that correspond to static obstacles. However, this occupancy map must differentiate between a static obstacle and a noisy LIDAR scan. We do this by checking consecutive LIDAR readings for obstacles that block grid cells in the occupancy grid and has not moved for some time. While this step increases the time between detecting an obstacle and replanning, it is necessary to prevent faulty replanning of the coverage path as a result of LIDAR scan noise/fluctuations.
Note that without replanning (i.e. GD replan with the local planner), the coverage path for the narrow regions will have short ranks with turns occurring in quick succession. The Neo robot cannot drive such paths without manual assistance. The proposed replanning approach aims to reduce such instances where manual assistance is required to continue coverage.
(a) Neo robot in Scenario 1
(b) Scenario 2 obstacles
Figure 17: Snapshots from experiments using the Avidbots Neo robot
(a) Initial Path
(b) Scenario 1
(c) Scenario 2
Figure 18: Robot experiment results of replanning an initial coverage path using OARP-Replan. The final robot path (pink curves) and the covered area (blue) are shown for both scenarios.
IX Conclusion
@@
We proposed an anytime coverage replanning approach called OARP-Replan to replan coverage paths for environments with unknown static obstacles. OARP-Replan solves this problem in two stages: (i) rank replanning, and (ii) touring. The first stage replans the ranks (straight-line paths) of an obstructed coverage path, while the second stage computes a tour of the replanned ranks to obtain the new path. We proposed two integer linear programs (ILPs) that each solve the rank replanning stage so that the resulting touring stage is completed within a time budget $\tau$. Simulations on maps of real-world environments showed that the OARP-Replan variants using each ILP achieve better coverage performance than a greedy replanner and state-of-the-art coverage planners used for online replanning [6, 18, 32]. Moreover, we also compared OARP-Replan with offline planning approaches (treating all obstacles as known) [7, 8] where we observed an improvement in total coverage time. Finally, we demonstrated the proposed approach by replanning coverage paths for an industrial-level cleaning robot.
@@
In this work, we focus only on the case of new obstacles in the environment. The case where an obstacle is observed and later removed from the environment (including dynamic obstacles) is left for future work. This would involve detecting areas in the environment that could be free at a later time due to an obstacle removal and returning to cover them without long detours. Another future direction involves replanning using predictions about environment occupancy to balance coverage and exploration, which are competing objectives. To balance them, the coverage planner must determine when to replan the coverage path, perhaps even delaying a replan until more information about an obstacle is collected. Removing the axis-parallel constraint and allowing multiple directions of coverage may also improve coverage paths for non-orthogonal environments.